OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is parallel to the line y = 3x +2 and passes through the point (5, 2). A. y = 3x – 5 B. y = 3x – 13 C. y = 3x – 1 D. y = 3x + 2

OpenStudy (lalaly):

y=mx+c parallel means they have the same slope... which means that m=3 y=2 , x=5 2=3(5)+c 2=15+c c=-13 so the equaation is y=3x-4

OpenStudy (lalaly):

oops y=3x-13 *

OpenStudy (anonymous):

thanks